### v1

# Microphysical Origins of Gravitational Potentials in Astrophysical Systems

Abstract
Most astrophysical systems (except for very compact objects such as, e.g., black holes and neutron stars) in our universe are characterized by shallow gravitational potentials, with dimensionless compactness $\left|\Phi\right|\equiv r_{s}/R\ll 1$, where $r_{s}$ and $R$ are their Schwarzschild radius and typical size, respectively. While the existence and characteristic scales of such virialized systems depend on gravity, we demonstrate that the value of $\Phi$—and thus the non-relativistic nature of most astrophysical objects—arises from microphysical parameters, specifically the fine structure constant and the electron-to-proton mass ratio, and is fundamentally independent of the gravitational constant $G$. It then follows that large-scale astrophysical velocities and escape velocities associated with virialized systems are determined by electromagnetic and atomic physics, not by gravitation, and that the compactness $\Phi$ is always set by microphysical scales — even for the most compact objects, such as neutron stars, where $\Phi$ is determined by quantities like the pion-to-proton mass ratio. This observation, largely overlooked in the literature, explains why the universe is not dominated by relativistic, compact objects and connects the low entropy of the observable universe to underlying microphysics. Our results emphasize the central but underappreciated role played by dimensionless microphysical constants in shaping the macroscopic gravitational landscape of the universe. In particular, we clarify that this independence of $\Phi$ from $G$ applies specifically to entire, virialized, naturally formed astrophysical systems — such as stars, galaxies, and planets — that have reached equilibrium through self-gravity and microphysical processes. In contrast, arbitrary non-virialized subsystems (e.g., a piece cut from a planet) do not exhibit this property.
I Introduction
The sensitivity of physical processes to the values of fundamental constants, such as the fine-structure constant $\alpha_{e}$ and the proton-to-electron mass ratio $\mu\equiv m_{p}/m_{e}$ — where $m_{p}$ and $m_{e}$ are the proton and electron masses, respectively — has been the subject of extensive theoretical and observational investigation. Variations in these constants impact not only particle physics, but also astrophysical processes, the formation of complex structures, and the potential habitability of the Universe.
Comprehensive reviews [1, 2] provide detailed accounts of how changes in $\alpha_{e}$ or $\mu$ modify atomic energy levels, nuclear reaction rates, and the epochs of recombination and primordial nucleosynthesis. For instance, even a modest increase in $\alpha_{e}$ accelerates atomic binding, shifting recombination to higher redshifts and potentially suppressing structure formation by altering the acoustic scale. Conversely, a lower $\alpha_{e}$ delays recombination and weakens atomic binding, with cascading effects on the cosmic microwave background (CMB) and baryon acoustic oscillations (BAO). Anthropic arguments have long explored the constraints imposed by these constants (e.g., [3, 4, 5]). Other aspects have been investigated in, e.g., [6, 7, 8, 9, 10].
These and other explorations of how fundamental constants impact astrophysics and cosmology motivate us to question their possible—if sometimes indirect or hidden—role in shaping the velocity fields associated with typical astrophysical systems. Remarkably, velocities observed in such systems never exceed a threshold of a few thousand km/sec. For example, the Earth’s velocity with respect to the Hubble frame is $\lesssim 400$ km/sec. The escape velocities from Venus, Earth, Mars, Saturn, and the Sun are approximately $6$, $11$, $5$, $35$, and $620$ km/sec, respectively. The solar system velocity relative to the Galactic center is 220220220220240240240240 km/sec, and typical infall velocities of merging clusters and subclusters have never been observed to exceed a few thousand km/sec. The corresponding gravitational wells are characterized by $\Phi+\beta^{2}\sim 0$, where $\Phi$ is the gravitational potential in speed of light-squared, $c^{2}$, units ($\Phi\equiv r_{s}/R$ is the “compactness” parameter where $r_{s}$ and $R$ are its Schwarzschild radius and actual size, respectively), and $\beta$ is the velocity in speed of light units.
This fact is somewhat puzzling, as a much more probable state would be a Universe teeming with black holes (BHs) and neutron stars (NSs), with $|\Phi|\lesssim 1$ and typical velocities approaching the speed of light. Such a Universe would be thermodynamically closer to the final heat death state, in which all BHs are evaporated via Hawking radiation and the far-future asymptotic de Sitter (dS) space is filled with blackbody radiation characterized by the corresponding temperature $T_{dS}\sim 10^{-32}$ eV. Regardless of its initial state, the penultimate state of the Universe is expected to be populated by BHs and NSs. On small scales, such a Universe would be much less homogeneous than the one we actually observe; higher-entropy gravitational systems are simply more clustered. The puzzle, then, is why our Universe is so regular and homogeneous ($|\Phi|\ll 1$). A key insight developed in this work is that, for entire, virialized, and naturally-formed astrophysical systems—such as stars, planets, and galaxies—the dimensionless $\Phi$ is fundamentally set by microphysical parameters (such as $\alpha_{e}$, $\mu$, and other dimensionless ratios), and is independent of the value of the gravitational constant $G$. This property does not apply to arbitrary subsystems (for example, a fragment cut from a planet), but holds for systems that have reached equilibrium through self-gravity and the relevant microphysical processes. Although the dependence of astrophysical masses and radii on fundamental constants is well explored, the general $G$-independence of $\Phi$ for virialized systems is not explicitly discussed in the literature, and we aim to highlight and clarify this point throughout the paper.
This paper is organized as follows. In Section II, we summarize a few basic results from the literature on the typical masses, radii, and temperatures of certain virialized astrophysical systems. In Section III, we highlight the non-observability of the gravitational constant in the absence of a fixed reference scale, and we summarize in Section IV. Throughout, we adopt natural units where the Boltzmann constant $k_{B}$, Planck constant $\hbar$, and the speed of light $c$ are all set to unity. Our estimates are only crude order-of-magnitude calculations, but this is sufficient for our purposes.
II Dimensional Analysis Considerations
Dimensional analysis is a powerful technique that generally allows for order-of-magnitude estimates of key quantities, encapsulating the main physical ingredients in a transparent fashion without recourse to differential equations (e.g., [11, 12, 13, 14, 15, 16, 17, 18, 19]). It captures the qualitative physics and underlying microphysics, leaving the details and precision to the full theory.
Quite generally, the average energy density, $\rho$, and temperature, $T$, characterizing an astrophysical system of mass $M$ and radius $R$, are given by
$M/R^{3}$ $\sim$ $\rho$ (1)
$GM/R$ $\sim$ $r_{s}/R\sim\Phi\sim T/m_{p}.$ (2)
If both $\rho$ and $T$ on the right-hand sides of Eqs. (1) and (2) are determined by the relevant microphysics of the system in question, then the system of Eqs. (1) and (2) can be solved for $R$, resulting in $R\sim\left(\frac{Tm_{p}}{\alpha_{g}\rho}\right)^{1/2}$ where $\alpha_{g}\equiv Gm_{p}^{2}$. It then follows that $M\sim\rho\left(\frac{Tm_{p}}{\alpha_{g}\rho}\right)^{3/2}$. Crucially, Eq. (2), $\Phi\sim T/m_{p}$, implies that the compactness $\Phi$—not only the density $\rho$—is completely decoupled from $G$. The latter only affects the typical values of $M$ and $R$. For example, assuming $\rho$ and $T$ are fully determined by miscrophysics and are independent of $G$, it immediately follows that $M\propto G^{-3/2}$ and $R\propto G^{-1/2}$. Exceptions are galaxies (and potentially galaxy clusters as well), whose typical $\rho$ does depend on $G$, and so $M$ and $R$ have different $G$-dependencies. Nevertheless, since even in these systems $T$ is independent of $G$, then $\Phi\sim T/m_{p}$ is independent of $G$, and the $G$-independence of the corresponding escape/virial velocities, gravitational lensing and gravitational redshift automatically follows.
The temperature $T$ is a property of the system. For example, the temperature of planets cannot exceed the Rydberg energy (in fact, it must be much lower; see [13]), the temperature of fuel burning stars cannot be smaller than $\sim 1$ keV, otherwise the Coulomb barrier cannot be overcome to allow for nuclear burning, etc. The idea is that temperature of virialized systems sets a lower limit on the escape velocity, and in case that the thermal velocity exceeds the escapse velocity the system quickly disperses and disintegrates. However, the escape velocity sets the gravitational potential, i.e., the ratio $M/R$. If in addition $\rho\sim M/R^{3}$ is known from the microphysics of the system in question, then as argued above, these two constraints can be solved for the characteristic $M$ and $R$.
Clearly, this conclusion is not inconsistent with the fact that $\Phi=GM/R$, and so for arbitrarily chosen $M$ and $R$, the gravitational potential $\Phi$ (and thus escape velocities and gravitational redshift) does depend on $G$. In other words, if we cut a piece of size $R_{0}$ from, e.g., a planet, its energy density does not change, but if $R_{0}$ (and consequently $M\sim\rho R_{0}^{3}$) is arbitrarily chosen, then $\Phi\sim G\rho R_{0}^{2}$ does depend on our arbitrary choice of $R_{0}$ as well as on $G$.
Two immediate consequences of all this are as follows. First, ’line broadening’ $\sigma_{\nu}=\sqrt{\frac{kT}{mc^{2}}}\nu_{0}\sim v\nu_{0}$ around a frequency $\nu_{0}$, caused by thermal motion of the emitters in a gravitationally bound system—essentially a redshift smearing effect—is independent of $G$. Second, although gravitational lensing is sourced by gravity, it is now trivial to see that the deflection angle is independent of $G$. This simply follows from the fact that the geodesic equations depend only on $\Phi$, but the latter is independent of $G$, and it then readily follows that the deflection angle is independent of $G$. $G$ could be a billion times larger or smaller than its measured value, yet the deflection angle would not change. More formally, the deflection angle $\alpha$ is determined by the transverse impulse, i.e., the ratio of the transverse photon momentum change to its momentum, $\alpha=2\int\nabla_{\perp}\Phi\,dl$. Now, since $\Phi$ is typically independent of $G$ in virialized astrophysical systems, then whatever $G$-dependence length scales might have, the combination $\alpha=2\int\nabla_{\perp}\Phi\,dl$ remains independent of $G$, as is appropriate for a geometric effect.
Let us inspect a few examples more closely. Our first example is rocky habitable planets. Following the arguments of [13], the temperature in habitable planets cannot exceed a certain threshold, specifically some fraction $\epsilon<1$ of the Rydberg energy, $Ry\equiv\frac{m_{e}}{2}\alpha_{e}^{2}$, so as not to disrupt the chemistry required for sustaining living organisms, i.e., $|\Phi|\sim v_{esc}^{2}\lesssim T/m_{p}$ where $T=Ry$. All this implies, e.g., [13], that $|\Phi|\sim\epsilon\alpha_{e}^{2}\mu^{-1}$. In other words, since $T<1$ eV and $m_{p}\lesssim 1$ GeV, then $|\Phi|=r_{s}/R\lesssim 10^{-9}$. For planet Earth, with radius $R\lesssim 10^{4}$ km, it implies that the corresponding Schwarzschild radius is $r_{s}\sim 1$ cm. The typical energy density of planets is given by $\rho\sim m_{p}/a_{0}^{3}$, where $a_{0}\equiv(m_{e}\alpha_{e})^{-1}$ is the Bohr radius, i.e., each Bohr volume is populated by approximately a single atom. Together, these simple atomic physics-based estimates of $T$ and $\rho$ uniquely determine $R$ and $M$ as described in Equations (1) and (2). In the case of planets in general, we require $T\sim Ry$ because the atomic structure of matter has to be preserved, albeit the chemistry of living organisms is no longer required to be protected. Thus, we require $T\sim Ry$ and so $\Phi\sim\alpha_{e}^{2}\mu^{-1}$ in this case, and $R$ and $M$ are correspondingly larger than in the habitable planets case by factors of $\epsilon^{-1/2}$ and $\epsilon^{-3/2}$, respectively.
Another example is nuclear-burning stars. Based on very general nuclear kinematic arguments a typical stellar temperature $T\sim 0.1\alpha_{e}^{2}m_{p}\sim 10$ keV is deduced in [14]. It then immediately follows that $|\Phi|\sim 0.1\alpha_{e}^{2}\sim 10^{-5}$. Here again, it is the weakness of the electromagnetic interaction that allows penetration through the Coulomb barrier and the ensuing nuclear fusion to take place. Had $\alpha_{e}$ been significantly larger, nuclear reactions would be impossible, resulting in matter collapsing under its own gravity into a black hole with $|\Phi|\sim 1$.
In [15], the estimate of galactic-scale systems required that for a gas cloud to fragment and form stars, the typical temperature has to drop below the Rydberg temperature. This by itself determines $|\Phi|\sim\alpha_{e}^{2}\mu^{-1}$. Another constraint used in [15] was that the dynamical gravitational time is comparable to the (Thomson) cooling time. Together, these constraints resulted in a fair estimate of $M$ and $R$. However, this early analysis ignored the role of CDM. It was revised later in [16] to account for CDM. As expected, the conclusion that $|\Phi|$ is shallow is robust against this improved analysis.
Molecular clouds do not seem to have been discussed in the literature in the context of typical scale estimates. Very general dimensional arguments imply that the vibrational and rotational molecular energy levels are given by $E_{vib}\sim\alpha_{e}^{2}m_{e}\mu^{-1/2}$ and $E_{rot}\sim\alpha_{e}^{2}m_{e}\mu^{-1}$. Considering virialized molecular clouds, the corresponding compactness parameters $\Phi$ are then $|\Phi|\sim\alpha_{e}^{2}\mu^{-3/2}$ and $|\Phi|\sim\alpha_{e}^{2}\mu^{-2}$, respectively, once again demonstrating the shallowness of $|\Phi|$. These estimates are general and decoupled from any assumptions pertaining to the corresponding $M$ and $R$.
White dwarfs (WD) and neutron stars (NS) are held against their own gravity by degeneracy pressure. In the case of WD, the degeneracy is that of electrons with typical spacing of order the Compton wavelength of electrons, naively $T\sim m_{e}$, and the corresponding $\Phi\sim\mu^{-1}$. In the case of NS, the effective temperature is $m_{\pi}$ and so $\Phi\sim m_{\pi}/m_{p}\lesssim 1$. Thus, only when gravity is opposed by the strong interaction, or by no opposing force at all, is $|\Phi|\sim 1$ and the corresponding escape velocities become relativistic.
These results, along with others based on references [13, 14, 15, 16] for various virialized astrophysical systems, are compiled and summarized in Table 1. The table presents typical masses, radii, number densities, and compactness in terms of the fine-structure constant $\alpha_{e}$, the proton-to-electron mass ratio $\eta\equiv m_{p}/m_{e}$, the proton-to-pion mass ratio $\kappa\equiv m_{p}/m_{\pi}$, the relative electromagnetic-to-gravitational interaction strength $\lambda\equiv\alpha_{e}/\alpha_{g}$ (where $\alpha_{g}\equiv Gm_{p}^{2}$), and a dimensionless parameter $\epsilon\sim 0.1$.
It is worth noting that, while the primary focus of the present work is the rightmost column showing the compactness parameter for a few typical astrophysical systems, the information in the other columns illustrates the valuable insights gained from basic dimensional analysis. For example, planets are less massive than their host stars fundamentally because $\alpha_{e}<1$, NS’s are less massive than their progenitor stars because $m_{\pi}<m_{p}$, and the observed Sun-to-Jupiter mass ratio, $O(10^{3})$, is simply determined by $\alpha_{e}^{-3/2}$. Likewise, the star-to-habitable planet (such as Earth) mass ratio is given by $\sim(\epsilon\alpha_{e})^{-3/2}$. While these observed ratios are typically explained by the details of various astrophysical processes, these processes are ultimately governed by dimensionless ratios of fundamental constants of nature, as demonstrated here.
$M/m_{p}$ $R/a_{0}$ $na_{0}^{3}$ $|\Phi|$
Habitable Planets $\epsilon^{\frac{3}{2}}\lambda^{\frac{3}{2}}$ $\epsilon^{\frac{1}{2}}\lambda^{\frac{1}{2}}$ $1$ $\sim\epsilon\alpha_{e}^{2}\mu^{-1}$
Planets $\lambda^{\frac{3}{2}}$ $\lambda^{\frac{1}{2}}$ $1$ $\sim\alpha_{e}^{2}\mu^{-1}$
Burning Stars $\alpha_{e}^{-\frac{3}{2}}\lambda^{\frac{3}{2}}$ $10\alpha_{e}^{-\frac{3}{2}}\lambda^{\frac{1}{2}}\eta^{-1}$ $(\alpha_{e}\mu/10)^{3}$ $\sim 0.1\alpha_{e}^{2}$
Galaxies $\lambda^{2}\alpha_{e}^{3}\mu^{\frac{1}{2}}$ $\alpha_{e}^{3}\mu^{\frac{1}{2}}\lambda$ $\alpha_{e}^{-6}\lambda^{-1}\mu^{-1}$ $\sim\alpha_{e}^{2}\mu^{-1}$
White Dwarfs $\alpha_{e}^{-\frac{3}{2}}\lambda^{\frac{3}{2}}$ $\alpha_{e}^{\frac{1}{2}}\lambda^{\frac{1}{2}}$ $\alpha_{e}^{-3}$ $\sim\mu^{-1}$
Neutron Stars $\alpha_{e}^{-\frac{3}{2}}\lambda^{\frac{3}{2}}[\kappa^{-3},\kappa]$ $\alpha_{e}^{\frac{1}{2}}\lambda^{\frac{1}{2}}\mu^{-1}$ $(\mu/\alpha_{e})^{3}[\kappa^{-3},\kappa]$ $[\kappa^{-3},\kappa$]
Table 1: A compilation based on [13, 14, 15, 16] of typical dimensionless values for various virialized astrophysical systems: typical masses in proton mass units $M/m_{p}$ (essentially $N$, the number of protons composing the system), radii in Bohr radius units $R/a_{0}$, number densities in Bohr units $na_{0}^{3}$ and the ‘compactness’ prameter, i.e. the gravitational potential in $c^{2}$ units, $|\Phi|$. Throughout, $\mu\equiv m_{p}/m_{e}$ and $\kappa\equiv m_{p}/m_{\pi}$ are the proton-to-electron and propton-to-pion mass ratios, respecively, $\lambda\equiv\alpha_{e}/\alpha_{g}$ is the electromagnetic-to-gravitational coupling strength ratio, and $\epsilon\sim 0.1$. In the case of NS a range of masses, densities and compactness are shown. The table highlights that except for the case of galaxies $na_{0}^{3}$ (or $n$ for that matter) are independent of $G$. The compactness parameters, $|\Phi|$, are independent of $G$, and in fact $|\Phi|\lesssim\alpha_{e}^{2}$ for various systems spanning a range of mass and length scales.
Thus, gravity, as encapsulated by the gravitational constant $G$, primarily determines the overall masses, radii, and particle numbers of gravitationally bound systems, setting the scale for the existence of macroscopic astrophysical objects. However, once such systems form and reach virial equilibrium, the depth of their gravitational wells, as characterized by the dimensionless compactness $\Phi$, is dictated by microphysical parameters and remains fundamentally independent of the value of $G$. In this sense, $G$ enables the presence of large-scale structures, but does not directly govern the characteristic velocities or compactness of these virialized systems.
III On the Observability of the Gravitational Constant $G$
As we saw, a subtle but significant aspect of gravitational physics is the fashion in which the gravitational constant $G$ enters observable phenomena. In many gravitational systems — particularly those that are virialized, such as stars, planets, and galaxies — quantities accessible to observation (e.g., gravitational redshift, line broadening, lensing deflection angles, escape velocities, and virial velocities) are determined by ratios or dimensionless combinations of parameters, wherein $G$ cancels out.
Similarly, in cosmology, the Friedmann equation in a spatially flat Universe can be cast in terms of dimensionless density parameters
$H^{2}(z)=H_{0}^{2}\left[\Omega_{r}(1+z)^{4}+\Omega_{m}(1+z)^{3}+\Omega_{\Lambda}\right],$ (3)
where $H(z)$ is the Hubble function, i.e. expansion rate of the Universe, $\Omega_{i}=\rho_{i}/\rho_{c}$, $\rho_{i}$ is the energy density of the i’th species, and $\rho_{c}=3H_{0}^{2}/(8\pi G)$ is the critical density. Here, $G$ is present only in the definition of the critical density; the expansion history, as determined solely by the ratios $\Omega_{i}$, is independent of $G$. Likewise, the comoving distance $d_{c}=\int_{0}^{z}\frac{dz^{\prime}}{H(z^{\prime})}$ is independent of $G$. All this implies that the angular diameter- and luminosity-distance are G-independent as well.
Only by combining cosmological measurements with an independent, non-gravitational determination of an absolute energy density—such as the CMB temperature (and thus the photon energy density)—can we directly constrain $G$. Without a direct measurement of the CMB temperature or the local baryon density, observations on cosmological scales would be insensitive to the empirical value of $G$.
It is only in the case of non-virialized systems, or in the direct interaction between two effectively pointlike masses (e.g., the Earth–Sun system), that $G$ appears explicitly in observable quantities, as in Newton gravitational law
$F=\frac{GMm}{r^{2}}.$ (4)
In these cases, the force or acceleration is directly proportional to $G$ and can be measured unambiguously.
This pattern suggests that the role of $G$ is, in a sense, emergent: it is hidden in systems governed by self-gravity and revealed only when comparing gravitational effects to non-gravitational standards or when considering pointlike interactions. In this view, $G$ serves as a conversion factor between mass/energy and spacetime curvature, but its absolute value is often inaccessible in the absence of a fixed external scale. This observation highlights a profound feature of gravitational physics, with possible implications for our understanding of gravity as a fundamental versus emergent interaction.
IV Summary
Typical astrophysical velocities are almost always non-relativistic, and the gravitational wells $\Phi$ characterizing these astrophysical systems are correspondingly shallow, i.e., $|\Phi|=r_{s}/R\ll 1$, where $r_{s}$ and $R$ are their typical Schwarzschild radius and actual size, respectively. According to the generalized Second Law of thermodynamics, this implies that the entropy budget of our Universe is much smaller than the holographic bound, $S_{dS}\sim 10^{122}$.
This can be neatly explained by the fact that astrophysical systems are either regulated by cooling processes (governed by the electromagnetic interaction) or are required not to allow certain electromagnetic processes to occur. In other words, these systems must remain below specific temperature thresholds, otherwise their virial velocities would exceed their escape velocities and the systems would rapidly disperse. Typical temperatures are set by atomic or molecular physics, with $T\sim m_{e}\alpha_{e}^{2}$, so that virialized objects are typically characterized by $|\Phi|\sim\alpha_{e}^{2}\mu^{-\beta}$, where $\mu=m_{p}/m_{e}$ and $\beta$ is positive.
Had $\alpha_{e}$ and $\mu$ been of order unity, the entire Universe would be teeming with BH’s and NS’s, with typical velocities approaching the speed of light, shortly after the Big Bang—regardless of its initial entropy state. So why is the Universe not predominantly populated by BH’s and NS’s ? Simply because $\alpha_{e}\ll 1$ and $\mu\gg 1$.
While the reasoning laid out in this work is straightforward, the somewhat surprising conclusion—that although the existence of astrophysical systems depends on $G$ and specifically on its empirical value, certain properties characterizing these systems (e.g., gravitational redshift, line broadening, escape velocities, and gravitational lensing) are independent of $G$—appears to be overlooked in the literature. Specifically, the fact that large-scale velocities are determined solely by microphysics—essentially by the electromagnetic interaction rather than by gravitation—is not sufficiently appreciated. The main objective of this work was to emphasize this rather surprising fact. Although order-of-magnitude estimates of $M$, $R$ and $T$ for various astrophysical systems from first principles are discussed in the literature, the facts that (a) $\Phi$ is small due to the weakness of the electromagnetic interaction and the small electron-to-proton mass ratio, and (b) $\Phi$ is independent of $G$, seem to be overlooked, or at least not sufficiently acknowledged. In particular, we stress that this $G$-independence of $\Phi$ applies specifically to entire, virialized, naturally-formed astrophysical systems—such as stars, planets, and galaxies—that have formed and equilibrated through self-gravity and the relevant microphysical processes. Arbitrary non-virialized subsystems, like a piece cut from a planet, do not exhibit this property.
Acknowledgments
This research was supported by a grant from the Joan and Irwin Jacobs donor-advised fund at the JCF (San Diego, CA).